All variables are real numbers - express in simplified form
square root of 14xy^2 multiplied by the square root of 3xy^3

Answer by fractalier(6550)

Okay from
sqrt(14xy^2) * sqrt(3xy^3) =
sqrt(42x^2y^5) =
sqrt(x^2y^4)*sqrt(42y) =
xy^2 * sqrt(42y)
that's as far as this one goes...